DAY 20-JULY-1999 Shift summary

Alexander Ledovskoy (ledovsko@fnpxsr)
Tue, 20 Jul 1999 15:57:48 -0500

Shift Summary for DAY shift, 20-JULY-99

On Shift: H. White, S. Ledovskoy

Hours of beam available at all intensities (>3.8E+12): 5.15h
Hours of beam available at nominal intensity: 3.78h
Hours of E832 data (special runs not included): 8.50h
Hours of special runs (laser scans, muon runs...): 0.00h

Accounting of lost time: 2-3 spills for kfrend

========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: 2 Short accesses during Main Injector power supply
repairs to get Peter Shanahan's wallet and oxygen
analyzer for Wah.

========================================================================
Overview of shift: E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13952 7:00 New shift crew takes over.

7:15 Send email to Theo about swics database files

82 8:07 ktevana errors (unknown crate from stream 5).
KFREND fixes it.

92 8:20 MU2-MU3 DYC is busy. KFREND fixes it

9:03 Noticed the beam intensity went down to 3.4E12.
Called MCR. They are working on it.

9:35 Checked on Alcohol temp. -4.8C

150 9:35 Main Injector power supply problems. No beam.
(No word on an estimate)

9:50 Got in touch with Theo. He said he would take care of
new swics database file (and he did).

10:05 Made controlled access to rescue Peter Shanahan's
wallet.

178 10:15 Csi 1354 has 31.5% RMS. Back to normal next spill.

221 Beam is back but horizontal position at SEED2 1.3
(should be -3.5). Called MCR to complain

228 Beam Horizontal alignment is back to normal.

11:35 Checked on Alcohol temp. -4.4C Swapped Blue Ice.

279 CsI 681 has RMS 33.6%. Back to normal next spill

304 ktevana errors (missing FFFF from cr 53, stream 5).
KFREND fixes it.

315 ktevana errors (missing FFFF from cr 53, stream 5).
KFREND fixes it.

14:23 No beam. MI power supply trip. Estimate 30 min
We go on brief access to get Oxigen analyzer for Wah.

15:45 Next shift crew is not here yet.

========================================================================
Follow up Notes on Previous Shift

o Contact Theo/Ashkan about sw_sdaq.dbs database. Done.

========================================================================
Items to Follow up on Next Shift

- Keep the door of the gas shack closed. Wild animals like to go
in there and cause unpredictable damages.